directions:solve problems using two unknowns and asystem of two equations.
one number is 7 more than another number. the sum of these two numbers is 101. what are the numbers?

Answer by fanism(39)

x-7=y --> x-y=7 -- eq1
x+y=101 -- eq2
therefore, let's take eq2-eq1
0 + 2y = 94
y = 47
therfore x = 101-47 = 54